### v1

# Data-Efficient Low-Complexity Acoustic Scene Classification in the DCASE 2024 Challenge

Abstract
This article describes the Data-Efficient Low-Complexity Acoustic Scene Classification Task in the DCASE 2024 Challenge and the corresponding baseline system. The task setup is a continuation of previous editions (2022 and 2023), which focused on recording device mismatches and low-complexity constraints. This year’s edition introduces an additional real-world problem: participants must develop data-efficient systems for five scenarios, which progressively limit the available training data. The provided baseline system is based on an efficient, factorized CNN architecture constructed from inverted residual blocks and uses Freq-MixStyle to tackle the device mismatch problem. The baseline system’s accuracy ranges from 42.40% on the smallest to 56.99% on the largest training set.
Index Terms—  DCASE Challenge, Acoustic Scene Classification, data-efficiency, low-complexity, multiple devices
1 Introduction
Acoustic Scene Classification (ASC) aims at detecting the environmental context in which audio was captured, based on a short excerpt [1]. The environmental context is given as a set of pre-defined acoustic scene classes such as Metro station, Urban park, or Public square. Since its inception, the ASC task has been an integral part of the DCASE Challenge. Each year’s edition focused on one or multiple challenging machine learning aspects in addition to the supervised classification task itself. These aspects include open-set classification [2], constraints on the model’s size and computational complexity [3, 4, 5], and generalization across different recording devices [6, 3]. These additional problems target the real-world applicability of ASC systems; for instance, the methods should be robust to diverse recording devices and sufficiently lightweight to be deployable on embedded devices. In the 2024 edition [^1] Task Description Page: https://dcase.community/challenge2024/task-data-efficient-low-complexity-acoustic-scene-classification of the ASC task, an additional challenging real-world aspect is addressed: the limited availability of training data. This setting intends to spark research on data-efficient learning methods capable of achieving high classification performance given only a small number of labeled acoustic scene examples for training.
Figure 1: Overview of Data-Efficient Low-Complexity Acoustic Scene Classification. The system must be trained on five datasets of varying sizes, it must generalize to unseen recording devices, and it is required to be lightweight enough for inference on an embedded device (ED).
Figure 1 shows an overview of the task setup. The ASC systems must be trained on subsets of a fixed training set that progressively limit the number of training samples, where the smallest subset only contains 5% of the audio snippets in the full training set (see Section  3.2). The training procedure is not limited in terms of complexity and may be executed on large GPU servers. However, aligned with real-world requirements, the system must be lightweight for inference such that it can be deployed on embedded devices (see Section  3.3). Additionally, the developed ASC system must be robust to unseen recording devices. To test this ability, the test set includes audio clips recorded by new devices that are not available in the training sets (see Section  3.2).
The remainder of this paper is structured as follows: Section 2 briefly discusses the role of low-complexity constraints and the device generalization problem in previous editions of the task. Section 3 introduces the setup for Data-Efficient Low-Complexity Acoustic Scene Classification in the DCASE 2024 Challenge; the baseline system is presented in Section 4. The outcome of the task and conclusions will be added when the challenge results are available.
2 Previous Editions
The low-complexity aspect has already been investigated in previous DCASE challenges and has undergone several refinements. In the 2020 [3] and 2021 [4] editions, systems were limited with respect to model size, allowing 500 kB and 128 kB for non-zero parameters, respectively. In the 2022 edition [5], the complexity constraint additionally included computational complexity, allowing a maximum of 30 MMACs (million multiply-accumulate operations), modeled after Cortex-M4 devices. The maximum number of parameters was 128K, with the variable type fixed to INT8. The 2023 edition took this one step further and included model size and computational complexity as part of the ranking metric, requiring participants to tune the system’s performance–complexity trade-off. In response to the low-complexity requirements, training techniques such as Sparsification [7], Pruning [8], Quantization [9], or Knowledge Distillation [10] have been extensively studied, and efficient factorized CNN architectures [11, 12, 13] have been designed.
Besides low-complexity techniques, substantial research has been conducted on the device mismatch problem. Efforts to improve device generalization involved suppressing device information via normalization [9] and domain adaptation [14], balancing the devices by changing the sampling distribution [15] and augmenting audio segments with device translators [9], Freq-MixStyle [10], and device impulse response augmentation [16].
3 Task Setup
While low complexity and generalization across different recording devices are well-studied topics, the specific aspect of interest in the 2024 edition is the limited availability of acoustic scene data for training. Specifically, participants are encouraged to develop data-efficient systems and study techniques that can alleviate the data scarcity problem, such as using extensive audio augmentation methods, pre-training models on general-purpose audio datasets, or novel, creative ideas that challenge participants may come up with.
3.1 Dataset
The task is based on the TAU Urban Acoustic Scenes 2022 Mobile dataset [6, 3], as used in the 2022 and 2023 editions of the task [5]. The dataset provides one-second audio snippets with a sampling rate of 44.1 kHz in single-channel, 24-bit format and consists of recordings from ten acoustic scenes: airport, indoor shopping mall, metro station, pedestrian street, public square, street with medium level of traffic, travelling by a tram, travelling by a bus, travelling by an underground metro, and urban park.
The audio has been recorded in multiple European cities with four recording devices in parallel. The primary device, referred to as device A, is a high-quality binaural device, while devices B, C and D are commonly available consumer devices. Additionally, 10 simulated devices (S1-S10) are created using audio from device A and a set of impulse responses from mobile devices. For details on the dataset creation and the exact distribution of devices, please refer to [3].
The dataset is split into a development set and an evaluation set. The development set, consisting of 64 hours of audio, contains 3 real devices (A, B, and C) and 6 simulated devices (S1–S6). The evaluation set comprises five unseen devices (D and S7-S10) and two unseen cities, in addition to devices and cities overlapping with the development set. The evaluation set is used to rank submissions and therefore comes without corresponding scene labels. Device and city information is not provided for recordings in the evaluation set.
3.2 Data-Efficient Evaluation
The development set used for the 2024 challenge is the same one as used in the previous two years and described above. It comes with the same pre-defined split into a development-train and a development-test partition. The development-train set contains six devices (A, B, C, S1-S3), leaving three unseen devices (S4-S6) for the development-test set to measure the device generalization performance.
For the evaluation of data efficiency, this year’s setup introduces five pre-defined subsets that progressively limit the available training data and contain 100%, 50%, 25%, 10%, and 5% of the recordings in the development-train set. The distribution of acoustic scenes, cities, and recording devices is kept similar across all subsets. The smaller subsets are fully included in the larger ones, corresponding to the idea of progressively collecting more data.
Participants are allowed to submit up to three different systems that may be specialized for the different training set sizes. Each system must be trained on all five subsets, and the performances on the development-test set must be reported. A system is considered to be the same if its architecture and design choices (such as building blocks, features, data augmentation techniques, decision-making, etc.) remain the same. However, basic hyperparameters like the number of update steps, learning rate, batch size, or regularization strength may vary for training on the different subsets.
All systems must be trained only on the respective subset and the explicitly allowed external resources. The allowed external resources include general-purpose audio datasets, such as AudioSet [17] or FSD50K [18], but no datasets specific to acoustic scenes.
The leaderboard ranking score is computed as follows. First, class-wise macro-averaged accuracies for all $P=5$ development-train subsets and all $N$ submissions are computed. The accuracy of the $n$-th submission on the $p$subset is denoted as $ACC_{n,p}$. The scores are then aggregated by choosing the best-performing system for each subset and averaging the resulting accuracies.
$\mathrm{score}:=\frac{1}{P}\sum_{p\in\{5,10,25,50,100\}}\max_{n\in\{1,...,N\}}ACC_{n,p}$ (1)
The outlined setup encourages research into the following scientific questions: how does the performance of systems vary with the number of available labeled training samples? how can systems be adapted to better cope with the limited availability of labeled training data? can pre-training of low-complexity models on general-purpose audio datasets effectively mitigate the need for larger amounts of acoustic scenes?
3.3 System Complexity Requirements
The system complexity is limited in terms of model size and MMACs. The maximum memory allowance for model parameters is 128 kB, with no requirement regarding the numerical representation. That is, participants can trade off the number of parameters and the numerical representation. For example, the memory limit translates to 128K parameters when using 8-bit quantization, or 32K parameters when using 32-bit precision. The computational complexity is limited to 30 MMACs for the inference on a one-second audio segment. These complexity limits are modeled after Cortex-M4 devices (e.g., STM32L496@80MHz or Arduino Nano 33@64MHz).
Figure 2: Baseline system performance on the development-test set for training on the five subsets (5%, 10%, 25%, 50%, 100%) of the development-train split. The three lines depict the performances on seen devices (A, B, C, S1, S2, S3), unseen devices (S4, S5, S6), and all devices. All results are averages over five independent trials.
4 Baseline System
The baseline system is a simplified version of the top-ranked system submitted to the 2023 edition [19] . It is based on a receptive-field-regularized, factorized CNN design. Audio input is resampled to 32 kHz and converted to mel spectrograms using a 4096-point FFT with a window size of 96 ms and a hop size of approximately 16 ms, followed by a mel transformation with a filterbank of 256 mel bins. The system is trained for 150 epochs using the AdamW [20] optimizer and a batch size of 256. Freq-MixStyle [10] is applied to tackle the device mismatch problem, and time rolling of the waveform and frequency masking are used to augment the training data. The baseline system requires 29.4 MMACs for the inference on a one-second audio clip. The memory required for the model parameters amounts to 122.3 kB, resulting from the 61,148 parameters used in 16-bit precision (float 16). The code and a detailed description of the baseline system are available online [^2] Source Code: https://github.com/CPJKU/dcase2024_task1_baseline/tree/main.
As Figure 2 shows, the baseline performance increases monotonically as the number of audio segments available for training increases. The macro-average accuracy ranges from 42.40% for the smallest training subset (5%) to 56.99% accuracy for the full set (100%). The plot also shows the performance on devices seen during training (A, B, C, S1, S2, S3) and devices only contained in the development-test split (S4, S5, S6). The large performance gap between these two lines emphasizes the need for more sophisticated device generalization methods.
5 ACKNOWLEDGMENT
The LIT AI Lab is supported by the Federal State of Upper Austria. Gerhard Widmer’s work is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant agreement No 101019375 (Whither Music?).